# Adam Wunderlich and Aric Sanders

### v1

@@ Abstract
One of the fundamental endeavors in radio frequency (RF) metrology is to measure the power of signals, where a common aim is to estimate the peak-to-average power ratio (PAPR), which quantifies the ratio of the maximum (peak) to the mean value. For a finite number of discrete-time samples of baseband in-phase and quadrature (I/Q) white Gaussian noise (WGN) that are independent and identically distributed with zero mean, we derive a closed-form, exact formula for mean PAPR that is well-approximated by the natural logarithm of the number of samples plus Euler’s constant. Additionally, we give related theoretical results for the mean crest factor. After comparing our main result to previously published approximate formulas, we examine how violations of the WGN assumptions in sampled I/Q data result in deviations from the expected value of PAPR. Finally, utilizing a measured RF I/Q acquisition, we illustrate how our formula for mean PAPR can be applied to spectral analysis with spectrograms to verify when measured RF emissions are WGN in a given frequency band.
@@
peak-to-average power ratio (PAPR), crest factor, peak detection, max-hold, Rayleigh distribution
@@ I Introduction
Modulated RF signals are typically measured, processed, and stored in an equivalent baseband representation, called the in-phase and quadrature (I/Q) signal representation [7]. Furthermore, the noise level in the absence of a signal is typically dominated by thermal noise in the I and Q components that can be modeled as a stationary, zero-mean, white Gaussian process [8, 9]. Therefore, the expected PAPR of stationary white Gaussian noise (WGN) in I/Q data is of high interest to differentiate thermal noise from other noise and signal types.
@@
The difference between PAPR assessment for continuous-time and discrete-time measurements is often ignored [10], because the Nyquist-Shannon sampling theorem implies that a band-limited signal may be perfectly reconstructed from samples if the sampling rate is greater than or equal to the Nyquist rate. However, real signals are not perfectly band-limited, and the analog-to-digital (A/D) and digital-to-analog (D/A) conversion process is non-ideal [11]. Consequently, the PAPR of a discrete-time sample is not identical to the PAPR for the corresponding continuous-time signal. Bounds on the difference between the peak of a continuous-time, band-limited signal and the peak value of a discrete-time, over-sampled measurement were studied in [12, 13].
@@
Extrema theory for continuous-time, stationary random processes is reviewed in [14, 15]. In particular, notable results concerning the expected number of local peaks above a specified level per unit time were developed for band-limited, continuous-time signals in [16, 17, 18]. This theory was later applied by Ochai and Imai [19] to derive accurate expressions for the PAPR distribution of band-limited, continuous-time orthogonal frequency-division multiplexing (OFDM) communication signals.
@@
The present work concerns the expected PAPR for discrete-time, sampled I/Q measurements of WGN, which is relevant to digital signal analyzer measurements. Namely, using extreme value theory [20] and the theory of order statistics [21], we develop highly accurate, closed-form formulas for the mean PAPR and mean CF for a finite collection of $n$ independent, identically distributed (i.i.d.) I/Q samples of stationary, zero-mean WGN. To our knowledge, the theoretical results given here and their connection to measurement have not been presented in the literature. For this problem, approximate formulas for the mean PAPR were presented in [3, 22]. Also, exact cumulative distribution functions (CDFs) of PAPR and CF for WGN I/Q samples were given in [19], but not analyzed further. Compared to these prior works, we present additional theoretical insights, context, and applications. Namely, we examine how non-ideal I/Q measurements of thermal noise can violate the WGN assumptions and impact the expected PAPR. Further, we show how the expected mean PAPR formula can be applied to spectrograms, which are widely used to visualize the frequency content of a nonstationary signal as it varies in time, to verify when measured RF emissions in a given frequency band consist of thermal noise.
@@ II Theoretical Results
Consider a baseband signal, $X$, with in-phase and quadrature signal components, $X_{I}$ and $X_{Q}$, respectively. Denote the amplitude (or envelope) of $X$ as $A=\sqrt{|X_{I}|^{2}+|X_{Q}|^{2}}$ and the instantaneous power (or squared amplitude) [^1] For generality, we avoid assuming physical units and the requisite scale factors necessary to convert between squared amplitude and power units. as $P=|X_{I}|^{2}+|X_{Q}|^{2}$. Unless otherwise stated, we assume that the I/Q components can be modeled as WGN. Specifically, $X_{I}$ and $X_{Q}$, are taken to be i.i.d. normal random variables with zero mean and variance $\sigma^{2}$, denoted $X_{I}\sim\mathcal{N}(0,\sigma^{2})$ and $X_{Q}\sim\mathcal{N}(0,\sigma^{2})$.
@@
We begin with a review of established distribution theory for PAPR and CF. To our knowledge, the subsequent theoretical results are new unless stated otherwise.
@@ II-A Distribution Theory
Since $X_{I}\sim\mathcal{N}(0,\sigma^{2})$ and $X_{Q}\sim\mathcal{N}(0,\sigma^{2})$, it follows that the amplitude, $A$, has a Rayleigh distribution with scale parameter, $\sigma$, and the instaneous power, $P$, has an exponential distribution with rate parameter, $\lambda=1/(2\sigma^{2})$ [23, 24]. Here, we indicate this with the notation $A\sim\text{Rayleigh}(\sigma)$ and $P\sim\text{Exp}(1/(2\sigma^{2}))$, where the quantity in parentheses denotes the parameter value.
For $x\geq 0$, the CDFs for $A$ and $P$ are $F_{A}(x)=1-\exp[-x^{2}/(2\sigma^{2})]$ and $F_{P}(x)=1-\exp[-\lambda x]=1-\exp[-x/(2\sigma^{2})]$, and the mean values are $\text{E}[A]=\sigma\sqrt{\pi/2}$ and $\text{E}[P]=1/\lambda=2\sigma^{2}$ [24]. Applying the CDF transformation theorem for an increasing function of a random variable [25, Theorem 2.1.3], it follows that for $x\geq 0$, the CDFs of $\widetilde{A}=A/(\sigma\sqrt{2})$ and $\widetilde{P}=P/(2\sigma^{2})$ are $F_{\widetilde{A}}(x)=1-\exp[-x^{2}]$ and $F_{\widetilde{P}}(x)=1-\exp[-x]$. In terms of the previously introduced distribution notation, $\widetilde{A}\sim\text{Rayleigh}(1)$ and $\widetilde{P}\sim\text{Exp}(1)$, i.e., $\sigma=1$ and $\lambda=1$ for the Rayleigh and exponential distributions, respectively.
Suppose we have $n$ i.i.d. measurements of $X_{I}\sim\mathcal{N}(0,\sigma^{2})$ and $X_{Q}\sim\mathcal{N}(0,\sigma^{2})$, yielding $n$ i.i.d. measurements, $\widetilde{P}_{1},\widetilde{P}_{2},\ldots,\widetilde{P}_{n}$, of $\widetilde{P}\sim\text{Exp}(1)$ and $n$ i.i.d. measurements, $\widetilde{A}_{1},\widetilde{A}_{2},\ldots,\widetilde{A}_{n}$, of $\widetilde{A}\sim\text{Rayleigh}(1)$. By a well-known formula for the CDF of the maximum of $n$ i.i.d. samples from a known CDF [21, Eq. 2.1.1], the CDFs of PAPR and CF are
@@
To express the PAPR PDF on a decibel (dB) scale, we apply the PDF transformation theorem [25, Theorem 2.1.5] with $_{\text{dB}}=10\log_{10}(\text{PAPR})$ PAPR ) to obtain
@@ II-B Asymptotic Distributions
From extreme value theory [20], it is known that the distribution for the maximum of an i.i.d. sample from an exponential or Rayleigh distribution asymptotically approaches a Gumbel distribution [20, Sec. 5.2] as the sample size increases. Specifically, letting $G(z)=\exp[-\exp(-z)]$ denote the standard Gumbel CDF, $\lim_{n\rightarrow\infty}F_{\text{PAPR}}(\ln(n)+z)=G(z)$ and $\lim_{n\rightarrow\infty}F_{\text{CF}}(\sqrt{\ln n}+z/(2\sqrt{\ln n}))=G(z)$ [20, p. 105-108]. Hence, for large $n$,
@@
The above CDFs take the form of a general Gumbel distribution, which has CDF $F(x;a,b)=\exp[-\exp(-(x-a)/b)]$ and mean $a+\gamma b$, where $\gamma=0.5772156649\ldots$ is Euler’s constant [20, p. 185].
@@ II-C Mean Peak-to-Average Power Ratio
First, we obtain an approximate asymptotic expression for the mean PAPR from (7) and the formula for the mean of the Gumbel distribution given in Sec. II-B. Namely, for large $n$,
@@
An exact closed-form formula for the mean PAPR is derived from established results on order statistics for the standard exponential distribution, which imply that the mean of the $r$th order statistic, $\widetilde{P}_{(r)}$, is [24, Eq. 19.12]
@@
Efficient computation of $H_{n}$ for large $n$ can be carried out with the asymptotic approximation [26, p. 480]
@@ II-D Mean Crest Factor
First, from (8) and the formula for the mean of a Gumbel distribution given in Sec. II-B, it follows that for large $n$, the mean CF is roughly
@@
The above approximation was found in [27, Eq. (59)] by other means in the context of narrow-band sea wave maxima.
We can also use (11) to derive an upper bound for the mean CF. Since $\text{CF}=\sqrt{\text{PAPR}}$, and because square root is a concave function, Jensen’s inequality [25, Thm. 4.7.7] implies that $\text{E}[\text{CF}]\leq\sqrt{\text{E}[\text{PAPR}]}$. Thus, (11) yields
@@
An exact closed-form expression for the mean CF can be obtained from known results for the mean order statistics of a Rayleigh distributed sample. Namely, starting with [24, Eq. 18.82] and simplifying yields
@@
Instead, we express the mean CF as an integral that can be accurately evaluated for large $n$ with standard numerical quadrature routines. The probability integral transformation [25, Thm. 2.1.10] implies that the random variable obtained by transforming CF with its CDF, $U=F_{\text{CF}}(\text{CF})$, follows a (standard) uniform distribution on the interval $[0,1]$. Thus, transforming a standard uniform random variable, $U$, with the CF quantile function (inverse CDF) yields $Q_{\text{CF}}(U)=\text{CF}$, and hence, $\text{E}[\text{CF}]=\text{E}[Q_{\text{CF}}(U)]$. Writing the last expectation in integral form with (4) gives
@@
(a) Mean CF Approximations (Sec. II-D)
(b) Mean PAPR Approximations (Sec. III)
Figure 1: Relative error in approximations for mean CF and mean PAPR with WGN I/Q data.
@@ III Comparison with Previously Published Approximate Formulas for Mean PAPR
We compare the exact result (11) presented in Sec. II-C to two approximate formulas from the engineering literature for the mean PAPR of WGN in I/Q data. The first formula, given in a book on microwave measurements [3, p. 588], states that for noise-like signals, the expected PAPR, expressed in linear units, is roughly
@@
An equipment vendor application note gives the following formula, expressed in linear units, for the average PAPR found from “a combination of analysis, approximation and experimentation” [22, p. 17-18]
@@ IV Examples of WGN Assumption Violations
Our theoretical findings are built upon assumptions about the nature of the noise and the operational space of the instruments recording it. Namely, our results require that the recorded samples of in-phase (I) and quadrature (Q) components are zero-mean WGN, implying that the I and Q components are
@@
Below, we provide examples illustrating how breakdowns in the above assumptions can occur in practical measurements of noise and impact the expected PAPR. Conversely, when measuring thermal noise, significant deviations from the expected mean PAPR for WGN provide evidence that I/Q measurements are non-ideal and violate one or more of the above assumptions.
@@ IV-B I/Q Imbalance
In RF transmitters and receivers, I/Q imbalance (or mismatch) results from gain and phase mismatches on the I and Q branches [31, Sec. 2.7]. Let $\Delta\phi$ be the phase mismatch between the I and Q components, and let $\Delta_{G}=|G_{I}-G_{Q}|/G_{Q}$ be the relative gain mismatch, where $G_{I}$ and $G_{Q}$ are the gains for the I and Q branches. Also, let $A=1-\Delta G/2$ and $B=1+\Delta G/2$. Using matrix multiplication, I/Q imbalance in an RF receiver can be modeled as [31, Eq. (2.131)]
@@
Monte Carlo simulations were performed to demonstrate the impacts of I/Q imbalance on the expected PAPR. For each sample size, I/Q samples were independently drawn from standard normal $\mathcal{N}(0,1)$ distributions and transformed with (22) for $(\Delta G,\Delta\phi)=(0.05,10^{\circ}),(0.1,15^{\circ})$, and $(0.2,20^{\circ})$. Additionally, the case of no I/Q imbalance, $(\Delta G,\Delta\phi)=(0,0)$, was simulated for baseline comparison. Mean PAPR estimated from $10^{4}$ independent Monte Carlo trials for each sample size is plotted in Fig. 3 (Top) together with the theoretical value (11) for ideal I/Q WGN. In all cases, the 95% confidence intervals were smaller than ±$\pm 0.022$ dB. From this plot, we observe that the expected PAPR increased as the I/Q imbalance became more severe. For example, in the $(\Delta G,\Delta\phi)=(0.2,20^{\circ})$ case with $n=10^{4}$ samples, the mean PAPR increased by roughly $0.8$ dB over the baseline case with no I/Q imbalance. Fig. 3 (Bottom) plots PDFs for the $n=10^{4}$ case estimated with the KDE method described in Sec. IV-A and compares them with the theoretical PDF for ideal I/Q WGN given by (6). For both plots in Fig. 3 in the ideal case with no I/Q imbalance, there was excellent agreement between the Monte Carlo results and theory.

### v3

@@ Abstract
One of the fundamental endeavors in radio frequency (RF) metrology is to measure the power of signals, where a common aim is to estimate the peak-to-average power ratio (PAPR), which quantifies the ratio of the maximum (peak) to the mean value. For a finite number of discrete-time samples of baseband in-phase and quadrature (I/Q) white Gaussian noise (WGN) that are independent and identically distributed with zero mean, we derive a closed-form, exact formula for mean PAPR that is well-approximated by the natural logarithm of the number of samples plus Euler’s constant. Additionally, we give related theoretical results for the mean crest factor (CF). After comparing our main result to previously published approximate formulas, we examine how violations of the WGN assumptions in sampled I/Q data result in deviations from the expected value of PAPR. Finally, utilizing a measured RF I/Q acquisition, we illustrate how our formula for mean PAPR can be applied to spectral analysis with spectrograms to verify when measured RF emissions are WGN in a given frequency band.
@@
Crest factor (CF), max hold, peak detection, peak-to-average power ratio (PAPR), Rayleigh distribution.
@@ I Introduction
Modulated RF signals are typically measured, processed, and stored in an equivalent baseband representation, called the in-phase and quadrature (I/Q) signal representation [7]. Furthermore, the noise level in the absence of a signal is typically dominated by thermal noise in the in-phase (I) and quadrature (Q) components that can be modeled as a stationary, zero-mean, white Gaussian process [8, 9]. Therefore, the expected PAPR of stationary white Gaussian noise (WGN) in I/Q data is of high interest to differentiate thermal noise from other noise and signal types.
@@
The difference between PAPR assessment for continuous-time and discrete-time measurements is often ignored [10], because the Nyquist-Shannon sampling theorem implies that a band-limited signal may be perfectly reconstructed from samples if the sampling rate is greater than or equal to the Nyquist rate. However, real signals are not perfectly band-limited, and the analog-to-digital (A/D) and digital-to-analog (D/A) conversion process is non-ideal [11]. Consequently, the PAPR of a discrete-time sample is not identical to the PAPR for the corresponding continuous-time signal. Bounds on the difference between the peak of a continuous-time, band-limited signal and the peak value of a discrete-time, over-sampled measurement were studied in [12] and [13].
@@
Extrema theory for continuous-time, stationary random processes is reviewed in [14] and [15]. In particular, notable results concerning the expected number of local peaks above a specified level per unit time were developed for band-limited, continuous-time signals in [16, 17]. This theory was later applied by Ochai and Imai [18] to derive accurate expressions for the PAPR distribution of band-limited, continuous-time orthogonal frequency-division multiplexing (OFDM) communication signals.
@@
The present work concerns the expected PAPR for discrete-time, sampled I/Q measurements of WGN, which is relevant to digital signal analyzer measurements. Namely, using extreme value theory [19] and the theory of order statistics [20], we develop highly accurate, closed-form formulas for the mean PAPR and mean CF for a finite collection of $n$ independent, identically distributed (i.i.d.) I/Q samples of stationary, zero-mean WGN. To our knowledge, the theoretical results given here and their connection to measurement have not been presented in the literature. For this problem, approximate formulas for the mean PAPR were presented in [3] and [21]. Also, exact cumulative distribution functions (CDFs) of PAPR and CF for WGN I/Q samples were given in [18], but not analyzed further. Compared to these prior works, we present additional theoretical insights, context, and applications. Namely, we examine how non-ideal I/Q measurements of thermal noise can violate the WGN assumptions and impact the expected PAPR. Further, we show how the expected mean PAPR formula can be applied to spectrograms, which are widely used to visualize the frequency content of a nonstationary signal as it varies in time, to verify when measured RF emissions in a given frequency band consist of thermal noise.
@@ II Theoretical Results
Consider a baseband signal, $X$, with I/Q signal components, $X_{I}$ and $X_{Q}$. Denote the amplitude (or envelope) of $X$ as $A=\sqrt{|X_{I}|^{2}+|X_{Q}|^{2}}$ and the instantaneous power (or squared amplitude) [^1] For generality, we avoid assuming physical units and the requisite scale factors necessary to convert between squared amplitude and power units. as $P=|X_{I}|^{2}+|X_{Q}|^{2}$. Unless otherwise stated, we assume that the I/Q components can be modeled as WGN. Specifically, $X_{I}$ and $X_{Q}$, are taken to be i.i.d. normal random variables with zero mean and variance $\sigma^{2}$, denoted $X_{I}\sim\mathcal{N}(0,\sigma^{2})$ and $X_{Q}\sim\mathcal{N}(0,\sigma^{2})$.
@@
We begin with a review of the established distribution theory for PAPR and CF. To the best of our knowledge, the subsequent theoretical results are new unless stated otherwise.
@@ II-A Distribution Theory
Since $X_{I}\sim\mathcal{N}(0,\sigma^{2})$ and $X_{Q}\sim\mathcal{N}(0,\sigma^{2})$, it follows that the amplitude, $A$, has a Rayleigh distribution with scale parameter, $\sigma$, and the instantaneous power, $P$, has an exponential distribution with rate parameter, $\lambda=1/(2\sigma^{2})$ [22, 23]. Here, we indicate this with the notation $A\sim\text{Rayleigh}(\sigma)$ and $P\sim\text{Exp}(1/(2\sigma^{2}))$, where the quantity in parentheses denotes the parameter value.
For $x\geq 0$, the CDFs for $A$ and $P$ are $F_{A}(x)=1-\exp[-x^{2}/(2\sigma^{2})]$ and $F_{P}(x)=1-\exp[-\lambda x]=1-\exp[-x/(2\sigma^{2})]$, and the mean values are $\text{E}[A]=\sigma\sqrt{\pi/2}$ and $\text{E}[P]=1/\lambda=2\sigma^{2}$ [23]. Applying the CDF transformation theorem for an increasing function of a random variable [24, Theorem 2.1.3], it follows that for $x\geq 0$, the CDFs of $\widetilde{A}=A/(\sigma\sqrt{2})$ and $\widetilde{P}=P/(2\sigma^{2})$ are $F_{\widetilde{A}}(x)=1-\exp[-x^{2}]$ and $F_{\widetilde{P}}(x)=1-\exp[-x]$. In terms of the previously introduced distribution notation, $\widetilde{A}\sim\text{Rayleigh}(1/\sqrt{2})$ and $\widetilde{P}\sim\text{Exp}(1)$, i.e., $\sigma=1/\sqrt{2}$ and $\lambda=1$ for the Rayleigh and exponential distributions, respectively.
Suppose we have $n$ i.i.d. measurements of $X_{I}\sim\mathcal{N}(0,\sigma^{2})$ and $X_{Q}\sim\mathcal{N}(0,\sigma^{2})$, yielding $n$ i.i.d. measurements, $\widetilde{P}_{1},\widetilde{P}_{2},\ldots,\widetilde{P}_{n}$, of $\widetilde{P}\sim\text{Exp}(1)$ and $n$ i.i.d. measurements, $\widetilde{A}_{1},\widetilde{A}_{2},\ldots,\widetilde{A}_{n}$, of $\widetilde{A}\sim\text{Rayleigh}(1/\sqrt{2})$. By a well-known formula for the CDF of the maximum of $n$ i.i.d. samples from a known CDF [20, Eq. 2.1.1], the CDFs of PAPR and CF are
@@
To express the PAPR PDF on a decibel (dB) scale, we apply the PDF transformation theorem [24, Theorem 2.1.5] with $_{\text{dB}}=10\log_{10}(\text{PAPR})$ PAPR ) to obtain
@@ II-B Asymptotic Distributions
From extreme value theory [19], it is known that the distribution for the maximum of an i.i.d. sample from an exponential or Rayleigh distribution asymptotically approaches a Gumbel distribution [19, Sec. 5.2] as the sample size increases. Specifically, letting $G(z)=\exp[-\exp(-z)]$ denote the standard Gumbel CDF, $\lim_{n\rightarrow\infty}F_{\text{PAPR}}(\ln(n)+z)=G(z)$ and $\lim_{n\rightarrow\infty}F_{\text{CF}}(\sqrt{\ln n}+z/(2\sqrt{\ln n}))=G(z)$ [19, p. 105-108]. Hence, for large $n$
@@
The above CDFs take the form of a general Gumbel distribution, which has CDF $F(x;a,b)=\exp[-\exp(-(x-a)/b)]$ and mean $a+\gamma b$, where $\gamma=0.5772156649\ldots$ is Euler’s constant [19, p. 185].
@@ II-C Mean Peak-to-Average Power Ratio
First, we obtain an approximate asymptotic expression for the mean PAPR from (7) and the formula for the mean of the Gumbel distribution. Namely, for large $n$
@@
An exact closed-form formula for the mean PAPR is derived from established results on order statistics for the standard exponential distribution, which imply that the mean of the $r$th order statistic, $\widetilde{P}_{(r)}$, is [23, Eq. 19.12]
@@
Efficient computation of $H_{n}$ for large $n$ can be carried out with the asymptotic approximation [25, p. 480]
@@ II-D Mean Crest Factor
First, from (8) and the formula for the mean of a Gumbel distribution, it follows that for large $n$, the mean CF is roughly
@@
The above approximation was found in [26, Eq. (59)] by other means in the context of narrow-band sea wave maxima.
We can also use (11) to derive an upper bound for the mean CF. Since $\text{CF}=\sqrt{\text{PAPR}}$, and because square root is a concave function, Jensen’s inequality [24, Thm. 4.7.7] implies that $\text{E}[\text{CF}]\leq\sqrt{\text{E}[\text{PAPR}]}$. Thus, (11) yields
@@
An exact closed-form expression for the mean CF can be obtained from known results for the mean order statistics of a Rayleigh distributed sample. Namely, starting with [23, Eq. 18.82] and simplifying yields
@@
Instead, we express the mean CF as an integral that can be accurately evaluated for large $n$ with standard numerical quadrature routines. The probability integral transformation [24, Th. 2.1.10] implies that the random variable obtained by transforming CF with its CDF, $U=F_{\text{CF}}(\text{CF})$, follows a standard uniform distribution on the interval $[0,1]$. Thus, transforming a standard uniform random variable, $U$, with the CF quantile function (inverse CDF) yields $Q_{\text{CF}}(U)=\text{CF}$, and hence, $\text{E}[\text{CF}]=\text{E}[Q_{\text{CF}}(U)]$. Writing the last expectation in integral form with (4) gives
@@
Figure 1: Relative error in approximations for mean CF and mean PAPR with WGN I/Q data. (a) Mean CF approximations (see Section II-D). (b) Mean PAPR approximations (see Section III).
@@ III Comparison with Previously Published Approximate Formulas for Mean PAPR
We compare the exact result (11) presented in Section II-C to two approximate formulas from the engineering literature for the mean PAPR of WGN in I/Q data. The first formula, given in a book on microwave measurements [3, p. 588], states that for noise-like signals, the expected PAPR, expressed in linear units, is roughly
@@
An equipment vendor application note gives the following formula, expressed in linear units, for the average PAPR found from “a combination of analysis, approximation and experimentation” [21, p. 17-18]
@@ IV Examples of WGN Assumption Violations
Our theoretical findings are built upon assumptions about the nature of the noise and the operational space of the instruments recording it. Namely, our results require that the recorded samples of I and Q components are zero-mean WGN, implying that the I and Q components satisfy the following conditions:
@@
Below, we provide examples illustrating how breakdowns in the above assumptions can occur in practical measurements of thermal noise and impact the expected PAPR. Conversely, when measuring thermal noise, significant deviations from the expected mean PAPR for WGN provide evidence that I/Q measurements are non-ideal and violate one or more of the above assumptions.
@@ IV-B I/Q Imbalance
In RF transmitters and receivers, I/Q imbalance (or mismatch) results from gain and phase mismatches on the I and Q branches [30, Sec. 2.7]. Let $\Delta\phi$ be the phase mismatch between the I and Q components, and let $\Delta_{G}=|G_{I}-G_{Q}|/G_{Q}$ be the relative gain mismatch, where $G_{I}$ and $G_{Q}$ are the gains for the I and Q branches. Also, let $A=1-\Delta G/2$ and $B=1+\Delta G/2$. Using matrix multiplication, I/Q imbalance in an RF receiver can be modeled as [30, Eq. (2.131)]
@@
Monte Carlo simulations were performed to demonstrate the impacts of I/Q imbalance on the expected PAPR. For each sample size, I/Q samples were independently drawn from standard normal $\mathcal{N}(0,1)$ distributions and transformed with (22) for $(\Delta G,\Delta\phi)=(0.05,10^{\circ}),(0.1,15^{\circ})$, and $(0.2,20^{\circ})$. In addition, the case of no I/Q imbalance, $(\Delta G,\Delta\phi)=(0,0)$, was simulated for baseline comparison. Mean PAPR estimated from $10^{4}$ independent Monte Carlo trials for each sample size is plotted in Fig. 3 (top) together with the theoretical value (11) for ideal I/Q WGN. In all cases, the 95% confidence intervals were smaller than ±$\pm 0.022$ dB. From this plot, we observe that the expected PAPR increased as the I/Q imbalance became more severe. For example, in the $(\Delta G,\Delta\phi)=(0.2,20^{\circ})$ case with $n=10^{4}$ samples, the mean PAPR increased by roughly $0.8$ dB over the baseline case with no I/Q imbalance. Fig. 3 (bottom) plots PDFs for the $n=10^{4}$ case estimated with the KDE method described in Section IV-A and compares them with the theoretical PDF for ideal I/Q WGN given by (6). For both plots in Fig. 3 in the ideal case with no I/Q imbalance, there was excellent agreement between the Monte Carlo results and theory.